# Growth rates of sequences governed by the squarefree properties of its translates

### v1

@@ 1. Introduction
Figure 1. Implications between various properties for infinite sets $A$. The dashed line indicates the incompatibility of $P$ and $\overline{P}_{\infty}$ in the infinite case. Note how $P$ stands in opposition to the other six properties, which are tied to each other by various implications.
@@
Figure 2. Implications and equivalences between various properties for finite sets $A$.
@@ 1.3. Squarefree sums
Similarly to Theorem 4, we can slightly lower this exponential growth:
Theorem 7 (Sub-exponential sequence).
There exists an absolute constant $C$ and an infinite sequence $A=\{1=a_{1}<a_{2}<\dots\}\subset\mathcal{SF}$ with squarefree sums such that $a_{j}\leq\exp(Cj/\log j)$ for all $j\geq 2$. In particular, $A$ has properties $Q$, $\overline{P}$, and $\overline{P}_{\infty}$ (but not $P$).
We prove this result in Section 4.3. In analogy with Theorem 4, in Section 4.4 we show that for large enough $j$ one can take $C$ to be any constant larger than $4$.
In the opposite direction, we have
Theorem 8.
Suppose one has an infinite sequence $A=\{a_{1}<a_{2}<\dots\}$ with squarefree sums. Then $a_{j}\gg j^{4/3}$ for all $j$. In particular, $A$ has natural density zero.
Proof.
For each prime $p$, the residue class $b\pmod{p^{2}}$ and $b\pmod{p^{2}}$ cannot both intersect $A$ for any $b$. Hence $A$ lies in at most $p^{2}/2$ such residue classes. By the large sieve (see Theorem 19) we conclude that for all $x,Q\geq 1$, the number of elements of $A$ less than or equal to $x$ is at most
$\frac{x+Q^{4}}{\sum_{q\leq Q}\mu^{2}(q)}.$
By setting $Q=x^{1/4}$ we thus see from (1.2) that this number is $O(x^{3/4})$, whence $a_{j}\gg j^{4/3}$. ∎
By instead choosing $Q=(x/3)^{1/4}$, the interested reader can verify with (1.2) that the implied constant in the lower bound on $a_{j}$ is equal to $\frac{3^{7/3}}{2^{4/3}\pi^{8/3}}+o(1)\approx 0.24$. Making use of the uniform lower bound $|\mathcal{SF}\cap[Q]|\geq 53Q/88$ from [rogers], it can even be checked that $a_{j}>0.24j^{4/3}$ holds for all $j\geq 1$. In the spirit of [shao], one would expect to be able to improve upon this lower bound if one could establish an inverse sieve conjecture, but such conjectures remain out of reach at present.
@@ 1.9. Acknowledgments
Most of the above questions on the possible growth rates of sequences with the various mentioned properties, were recently added to Bloom’s website [EP, Problems 1102, 1103]. There they caught the authors’ attention, and as far as we are aware, no other work has been done on any of these questions in the intervening years since they were first posed.
@@ 4.1. The lower bound
and let $x$ be sufficiently large depending on $\varepsilon$ and $n$. Let $n^{\prime}$ be chosen uniformly at random among the multiples of $W$ in $[x/2,x]$. It will suffice to show that with probability 1o(1)1-o(1) (as $n\to\infty$), that $n^{\prime}$ obeys the two properties (a), (b) of the proposition.
We begin with (a). If $a\in\mathcal{SF}$, then for any prime $p\leq n^{2}$, $p^{2}$ divides $W$ and hence $n^{\prime}$, but not $n^{\prime}+a$. Also, if $a\leq n$, then $n^{\prime}+a\leq 2x$. Thus, if $n^{\prime}+a\notin\mathcal{SF}$, then $n^{\prime}+a$ must be divisible by $p^{2}$ for some $n^{2}<p\leq\sqrt{2x}$. On the other hand, from Lemma 12 we see that for each such $p$ and any $a\leq n$, the probability that $n+a$ is divisible by $p^{2}$ is $O(\frac{1}{p^{2}}+\frac{W}{x})$. By the union bound, the probability that (a) fails is then at most
@@
thanks to (2.1) and the assumption that $x$ is sufficiently large. The claim (a) follows.
Now we turn to (b) and we aim to show that the inequality
@@
thanks to (2.1). Summing over $n\leq R\leq x$ that are powers of $1+\varepsilon$, the total failure probability is then
@@
since $x$ is large, giving the claim. This concludes the proof of Proposition 14 and hence Theorem 3.
@@ 4.2. Sufficiently rapidly growing sequences
4.3. A sub-exponential sequence
We modify the previous arguments to prove Theorem 7. Let $c>0$ be small, let $C$ be sufficiently large depending on $c$, and write $x_{j}\coloneqq\exp(Cj/\log j)$ for all $j\geq 2$, so that $j\leq c\log x_{j}\log\log x_{j}$. With $a_{0}=0$ and $a_{1}=1$, we will recursively construct $a_{j}\in[x_{j}-x_{j}^{4/5},x_{j}]$ for $j\geq 2$ such that $a_{i}+a_{j}\in\mathcal{SF}$ for all $i$ with $0\leq i\leq j$. For $C$ large enough, the sequence $A=\{a_{1},a_{2},\dots\}$ will then be increasing and obey all the required properties.
Suppose that we have already constructed $a_{1},a_{2},\dots,a_{j-1}$ obeying the required properties. We set
$W_{j}\coloneqq\prod_{p\leq\frac{1}{10}\log x_{j}}p^{2}=x_{j}^{1/5+o(1)}$
as before, where $o(1)$ is with respect to the limit $x_{j}\to\infty$. We select $a_{j}$ uniformly at random from the elements of the residue class $1\pmod{W_{j}}$ in $x_{j}-x_{j}^{4/5},x_{j}]$, of which there are $\asymp x_{j}^{4/5}/W_{j}$ many. We claim that $a_{i}+a_{j}$ is not divisible by $p^{2}$ for any $p\leq\frac{1}{10}\log x_{j}$ and any $0\leq i\leq j$. For $i=j$ this is clear by the fact that $2a_{j}\equiv 2\pmod{W_{j}}$, while for $i<j$ one can see this by the fact that
$a_{i}+a_{j}\equiv a_{i}+1\equiv a_{i}+a_{1}\pmod{W_{j}}$,
while, by the recursive construction, $a_{i}+a_{1}\in\mathcal{SF}$ and hence not divisible by $p^{2}$. On the other hand, $a_{i}+a_{j}\leq 2x_{j}$ is not divisible by $p^{2}$ for any $p>\sqrt{2x_{j}}$ either. For each $\frac{1}{10}\log x_{j}<p\leq\sqrt{2x_{j}}$, the probability that any sum is divisible by $p^{2}$ is $O(1/p^{2}+W_{j}/x_{j}^{4/5})$, thanks to Lemma 12. Thus the total probability that $a_{i}+a_{j}\notin\mathcal{SF}$ for some $0\leq i\leq j$ is
$\ll(j+1)\sum_{\frac{1}{10}\log x_{j}<p\leq\sqrt{2x_{j}}}\left(\frac{1}{p^{2}}+\frac{W_{j}}{x_{j}^{4/5}}\right)$
which we can bound using (2.1) as before by
$\ll c\log x_{j}\log\log x_{j}\left(\frac{1}{\log x_{j}\log\log x_{j}}+\frac{x_{j}^{7/10+o(1)}}{x_{j}^{4/5}}\right)\ll c.$
For $c$ small enough, this is smaller than $1$, so that an $a_{j}$ with the required properties is seen to exist.
Remark 15.
With $kA$ defined as the set of sums of $k$ (not necessarily distinct) elements of $A$, the above proof shows the existence of a sequence $A$ for which both $A$ and $2A$ are contained in $\mathcal{SF}$. One can modify the proof to arrange for $3A$ to be contained in $\mathcal{SF}$ as well, by requiring $a_{i}+a_{i}^{\prime}+a_{j}\in\mathcal{SF}$ for all $0\leq i\leq i^{\prime}\leq j$. In order to do this, one needs to change $x_{j}$ to $\exp(Cj^{2}/\log j)$ (thereby increasing the growth rate of $A$), $\sqrt{2x_{j}}$ to $\sqrt{3x_{j}}$, and $j+1=\binom{j+1}{1}$ in the calculation of the probability that some sum is not contained in $\mathcal{SF}$ to $\binom{j+1}{2}$. We note that $A\cup 2A\cup 3A\subset\mathcal{SF}$ is a natural limit, as $4A$ will contain an element divisible by $4$, for all non-empty sets $A$. On the other hand, a further natural generalization gives, for every finite set $B=\{b_{1}<b_{2}<\ldots<b_{m}\}\subset\mathcal{SF}$ with $b_{m}\geq 2$, a sequence $A=\{a_{1}<a_{2}<\ldots\}$ with $a_{j}<\exp(Cj^{b_{m}-1}/\log j)$ such that $b_{i}A\subset\mathcal{SF}$ for all $1\leq i\leq m$.
4.4. Explicit versions
In Section 1 we claimed that one can take $C$ in both Theorem 4 and Theorem 7 to be any constant larger than $4$. To see this, we first note that we can improve (2.1) and the upper bound probability in Lemma 12 to $\sum_{p\geq x}\frac{1}{p^{2}}=\frac{1+o(1)}{x\log x}$ and $\frac{1}{q}+O(\frac{W}{L})$ respectively. Secondly, we remark that it is possible to increase the constant $\frac{1}{10}$ that appears in Section 4.2 and Section 4.3 in the definitions of $W$ and $W_{j}$, to $\frac{1}{4}-\epsilon$. For this, we then need to slightly increase the interval from which we choose the $a_{j}$ in the proof of Theorem 7, to $x_{j}-x_{j}^{1-O(\epsilon)},x_{j}]$. With $c=\frac{1}{4}-O(\epsilon)$ and $C=\frac{1}{c}+O(\epsilon)$, one can then check that the final calculations to upper bound the failure probabilities (which currently give $\ll c$ in both proofs) instead give quantities that are, for sufficiently large $x$ and $x_{j}$, smaller than $1$.

### v2

# Growth rates of sequences governed by the squarefree properties of their translates

@@ 1. Introduction
Figure 1. Implications between various properties for infinite sets AA. The dashed line indicates the incompatibility of $P$ and $\overline{P}_{\infty}$ in the infinite case. Note how $P$ stands in opposition to the other six properties, which are tied to each other by various implications.
@@
Figure 2. Implications and equivalences between various properties for finite sets AA.
@@ 1.3. Squarefree sums
Let $\mathrm{ES}_{N}$ denote the size of the largest subset of $[N]$ with squarefree sums. Erdős and Sárközy [erdos-sarkozy], Sárközy [sarkozy], Gyarmati [gyarmati] and Konyagin [konyagin] have given upper and lower bounds on this quantity, with the current best bounds of
$\log^{2}N\log\log N\ll\mathrm{ES}_{N}\ll N^{11/15}\exp\left(O\left(\frac{\log N}{\sqrt{\log\log N}}\right)\right)$
for all large $N$ established in [konyagin]. (For comparison, the large sieve inequality in Theorem 16 only gives the weaker upper bound $\mathrm{ES}_{N}\ll N^{3/4}$.) Inverting the upper bound, we obtain a lower bound
$a_{j}\gg j^{15/11}\exp\left(-O\left(\frac{\log j}{\sqrt{\log\log j}}\right)\right)$
for all large $j$ and any infinite sequence $A=\{a_{1}<a_{2}<\dots\}$ with squarefree sums. It is likely that the constructions in [konyagin] can be modified to generate an infinite sequence $A=\{a_{1}<a_{2}<\dots\}$ with squarefree sums such that
$\log^{2}N\log\log N\ll|A\cap[N]|$
for all large $N$, although we will not do so here. Assuming that this is the case, such a sequence would obey the upper bound
$a_{j}\ll\exp(O(j^{1/2}/\log^{1/2}j)).$
We remark that a modification of the construction in Theorem 4 gives an infinite sequence $A=\{a_{1}<a_{2}<\dots\}$ for which the weaker upper bound
$a_{j}\ll\exp(O(j/\log j))$
holds. We refer the reader to the first version of this paper [^3] https://arxiv.org/abs/2512.01087v1 for details.
We also note that the opposite problem of studying sets with very few (or no) squarefree pairwise sums was considered in [nathanson], [filaseta], [lacampagne], [schoen].
@@ 1.9. Acknowledgments
We thank Thomas Bloom for his website [EP] where the above problems were brought to the authors’ attention at [EP, Problems 1102, 1103]. We also thank Benjamin Bedert for references concerning sets with squarefree sums.
@@ 4.1. The lower bound
and let $x$ be sufficiently large depending on $\varepsilon$ and $n$. Let $n^{\prime}$ be chosen uniformly at random among the multiples of $W$ in $[x/2,x]$. It will suffice to show that $n^{\prime}$ obeys properties (i) and (ii) of the proposition with probability 1o(1)1-o(1) as $n\to\infty$.
We begin with (i). If $a\in\mathcal{SF}$, then for any prime $p\leq n^{2}$, $p^{2}$ divides $W$ and hence $n^{\prime}$, but not $n^{\prime}+a$. Also, if $a\leq n$, then $n^{\prime}+a\leq 2x$. Thus, if $n^{\prime}+a\notin\mathcal{SF}$, then $n^{\prime}+a$ must be divisible by $p^{2}$ for some $n^{2}<p\leq\sqrt{2x}$. On the other hand, from Lemma 10 we see that for each such $p$ and any $a\leq n$, the probability that $n+a$ is divisible by $p^{2}$ is $O(\frac{1}{p^{2}}+\frac{W}{x})$. By the union bound, the probability that (i) fails is then at most
@@
thanks to (2.1) and the assumption that $x$ is sufficiently large. The claim (i) follows.
Now we turn to (ii) and we aim to show that the inequality
@@
thanks to (2.1). Summing over $n\leq R\leq x$ that are powers of $1+\varepsilon$, the probability that (ii) fails is then
@@
since $x$ is large. This concludes the proof of Proposition 12 and hence Theorem 3.
@@ 4.2. Sufficiently rapidly growing sequences
4.3. Explicit version
In Section 1 we claimed that one can take $C$ in Theorem 4 to be any constant larger than $4$. To see this, we first note that we can improve (2.1) and the upper bound probability in Lemma 10 to $\sum_{p\geq x}\frac{1}{p^{2}}=\frac{1+o(1)}{x\log x}$ and $\frac{1}{q}+O(\frac{W}{L})$ respectively. Secondly, we remark that it is possible to increase the constant $\frac{1}{10}$ that appears in the definition of $W$, to $\frac{1}{4}-\epsilon$. With $c=\frac{1}{4}-O(\epsilon)$ and $C=\frac{1}{c}+O(\epsilon)$, one can then check that the final calculation to upper bound the failure probability (which currently gives $\ll c$) instead gives a quantity that is, for sufficiently large $x$, smaller than $1$.